A car of mass 1100 kg drives with a velocity such that it has a kinetic energy of 400 kJ. Find its speed. If the car is raised with a crane, how high should it be lifted in the standard gravitational field to have a potential energy that equals the kinetic energy?
1
Expert's answer
2017-02-21T08:20:22-0500
E = mv^2/2

hence we can find velocity v = sqrt(2E/m) = sqrt(2*400*10^3 J/1100) = 27 m/s

When potential energy equals kinetic

mgh=mv^2/2=400kJ

The height is

h = mv^2/(2m*g) = 400kJ/(1100kg*9.8 m/s^2) = 37.1 m